# Understanding the Training and Generalization of Pretrained Transformer for Sequential Decision Making

### v1

@@ D.4 Examples for the optimal decision function $\texttt{Alg}^{*}$ in Section 4
(2). For the dynamic pricing problem, we consider the following toy model. Consider two environments $\left\{\gamma_{1},\gamma_{2}\right\}$, the prior distribution $\mathcal{P}_{\gamma}$ is given by $\mathcal{P}(\gamma_{1})=\mathcal{P}(\gamma_{2})=\frac{1}{2}$. There is no context, and the action space is 1-dimensional. The demand functions of them are set to be $o_{t}=2-a_{t}+\epsilon_{t}$ and $o_{t}=\frac{4}{5}-\frac{1}{5}\cdot a_{t}+\epsilon_{t}$ respectively, where $\epsilon_{t}\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$. Accordingly, the optimal actions (which follow the same notations as in (1)) are then $a_{(\gamma_{1})}=1$ and $a_{(\gamma_{2})}=2$.
Given $\gamma$ and $H_{t}$, $a_{t}$ is achieved by the posterior averaging:
$a_{t}=\texttt{Alg}^{*}(H_{t})=\sum_{i=1}^{2}\mathcal{P}(\gamma_{i}|H_{t})\cdot a_{(\gamma_{i})}$
Condition on $\gamma,a_{1},o_{1},\ldots,a_{t}$, the conditional distribution of the demand $o_{t}$ is the normal distribution $\mathcal{N}\left(2-a_{t},1\right)$ for $\gamma_{1}$ and $\mathcal{N}\left(\frac{4-a_{t}}{5},1\right)$ for $\gamma_{2}$. Observe that $\tilde{a}=\frac{1}{2}a_{(\gamma_{1})}+\frac{1}{2}a_{(\gamma_{2})}=\frac{3}{2}$ satisfies that $2-\tilde{a}=\frac{4-\tilde{a}}{5}$ and $a_{1}=\tilde{a}$, then by the similar procedure in (1) we can conclude that $\mathcal{P}\left(\gamma_{i}|H_{t}\right)=\frac{1}{2}$ for any $t$. Therefore, $a_{t}=\frac{1}{2}a_{(\gamma_{1})}+\frac{1}{2}a_{(\gamma_{2})}=\frac{3}{2}$ does not change with respect to $t$. Therefore the regret for $\gamma$ is
@@
$\text{Regret}(\texttt{Alg}^{*};\gamma)=\mathbb{E}\left[\sum_{t=1}^{T}r(X_{t},a_{t}^{*})-r(X_{t},a_{t})\right]=\frac{1}{4}T\cdot\mathbb{I}\left\{\gamma=\gamma_{1}\right\}+\frac{1}{20}T\cdot\mathbb{I}\left\{\gamma=\gamma_{2}\right\},$
which completes the second part of the proof.
@@
In the following proof, we assume Assumption B.1 holds. Recall that the posterior condition on history $H_{t}$ is $\mathcal{P}(\gamma_{i}|H_{t})=\frac{\prod_{{\tau}=1}^{t-1}\mathbb{P}_{\gamma_{i}}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}{\sum_{i=1}^{k}\prod_{{\tau}=1}^{t-1}\mathbb{P}_{\gamma_{i}}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}$, then for $\gamma\neq\gamma_{i}$, the log-likelihood ratio is
$\log\left(\frac{\mathcal{P}(\gamma|H_{t})}{\mathcal{P}(\gamma_{i}|H_{t})}\right)=\sum_{{\tau}=1}^{t-1}\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}{\mathbb{P}_{\gamma_{i}}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}\right).$
Since $\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}{\mathbb{P}_{\gamma_{i}}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}\right)$ is $\sigma^{2}$-subgaussian condition on $H_{t}$ with respect to environment $\gamma$, and $\left\{\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}{\mathbb{P}_{\gamma_{i}}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}\right)-\mathbb{E}\left[\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}{\mathbb{P}_{\gamma_{i}}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}\right)\Big{|}H_{\tau}\right]\right\}_{{\tau}=1}^{t}$ is a martingale difference sequence, by the concentration bounds for martingale difference sequence (Section 2.2.2 in [Wainwright, 2019]), with probability at least $1-h$, we have
$\log\left(\frac{\mathcal{P}(\gamma|H_{t})}{\mathcal{P}(\gamma_{i}|H_{t})}\right)$ $\geq\sum_{{\tau}=1}^{t-1}\mathbb{E}\left[\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}{\mathbb{P}_{\gamma_{i}}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}\right)\Big{|}H_{\tau}\right]-\sqrt{2\sigma^{2}(t-1)\log\left(\frac{1}{h}\right)}$
$\geq(t-1)\Delta-\sqrt{2\sigma^{2}(t-1)\log\left(\frac{1}{h}\right)}.$
[…]
@@
$\log\left(\frac{\mathcal{P}(\gamma|H_{t})}{\mathcal{P}(\gamma_{i}|H_{t})}\right)$ $\geq(t-1)\Delta-\sqrt{2\sigma^{2}(t-1)\log\left(\frac{k}{h}\right)},$
which implies that
$\mathcal{P}(\gamma|H_{t})\geq\exp\left((t-1)\Delta-\sqrt{2\sigma^{2}(t-1)\log\left(\frac{k}{h}\right)}\right)\mathcal{P}(\gamma_{i}|H_{t}).$
In view of this, we have
$0\leq\sum_{\gamma_{i}\neq\gamma}\mathcal{P}(\gamma_{i}|H_{t})\leq\frac{k-1}{k-1+\exp\left((t-1)\Delta-\sqrt{2\sigma^{2}(t-1)\log\left(\frac{k}{h}\right)}\right)}.$
Notice that the term $k-1$ in the denominator, derived from the condition $\mathop{\sum}\limits_{i=1}^{k}\mathcal{P}(\gamma_{i}|H_{t})=1$, is essential for getting the final regret bound.
Now we start to bound the regret. Recall that under the square loss, $\texttt{Alg}^{*}(H_{t})=\mathop{\sum}\limits_{i=1}^{k}a^{*}(H_{t},\gamma_{i})\mathcal{P}(\gamma_{i}|H_{t})$, where $a^{*}(H_{t},\gamma_{i})$ is the optimal action in environment $\gamma_{i}$. Then conditioned on $H_{t}$, with respect to $\gamma$ we have
$\mathbb{E}\left[r(X_{t},a_{t}^{*})-r(X_{t},\texttt{TF}_{\hat{\theta}}(H_{t}))|H_{t}\right]$
$=$ $\mathbb{E}\left[r(X_{t},a^{*}(H_{t},\gamma))-r\left(X_{t},\texttt{Alg}^{*}(H_{t})\right)|H_{t}\right]+\mathbb{E}\left[r\left(X_{t},\texttt{Alg}^{*}(H_{t})\right)-r\left(X_{t},\texttt{TF}_{\hat{\theta}}(H_{t})\right)|H_{t}\right].$
Recall that from Proposition B.2, the regret satisfies the surrogate property with respect to squared loss
$r(X_{t},a^{*}(H_{t},\gamma))-r\left(X_{t},\texttt{Alg}^{*}(H_{t})\right)\leq C||a^{*}(H_{t},\gamma)-\texttt{Alg}^{*}(H_{t})||_{2}^{2},$
this yields
$\mathbb{E}\left[r(X_{t},a_{t}^{*})-r(X_{t},\texttt{TF}_{\hat{\theta}}(H_{t}))|H_{t}\right]$
$\leq$ $C\mathbb{E}\left[||a^{*}(H_{t},\gamma)-\texttt{Alg}^{*}(H_{t})||_{2}^{2}\mid H_{t}\right]+\epsilon$
$=$ $C\mathbb{E}\left[\Big{|}\Big{|}\sum_{\gamma_{i}\neq\gamma}^{m}\left(a^{*}(H_{t},\gamma)-a^{*}(H_{t},\gamma_{i})\right)\mathcal{P}(\gamma_{i}|H_{t})\Big{|}\Big{|}_{2}^{2}\Big{|}H_{t}\right]+\epsilon$
$\leq$ $CD^{2}h+CD^{2}(1-h)\frac{k-1}{k-1+\exp\left((t-1)\Delta-\sqrt{2\sigma^{2}(t-1)\log\left(\frac{k}{h}\right)}\right)}+\epsilon,$
where $D$ is the diameter of the action space. Hence we have
$\mathrm{Regret}(\texttt{TF}_{\hat{\theta}};\gamma)=$ $\mathbb{E}\left[\sum_{t=1}^{T}r(X_{t},a^{*}(H_{t},\gamma))-r\left(X_{t},\texttt{TF}_{\hat{\theta}}(H_{t})\right)\right]$
$\leq$ $CD^{2}\left(hT+\sum_{t=1}^{T}\frac{k-1}{k-1+\exp\left((t-1)\Delta-\sqrt{2\sigma^{2}(t-1)\log\left(\frac{k}{h}\right)}\right)}\right)+\epsilon T$
$\leq$ $O\left((\epsilon+h)T+\frac{\log k}{\Delta}-\frac{\log h}{\Delta}\right).$
Taking $h=\epsilon$ completes the proof.
D.6.1 Justification for the conditions in Proposition 4.4
From now on, we provide some justification for the conditions in Proposition 4.4.
(1). We illustrate two examples (linear badits and dynamic pricing) to show the validity of the condition that the conditional log-likelihood ratio $\log\left(\frac{\mathbb{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})}{\mathbb{P}_{\gamma_{j}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})}\right)$ is $\sigma^{2}$-subgaussian with respect to $\gamma_{i}$.
  • For linear bandits problem, given $\gamma,a_{1},o_{1},\ldots,a_{t}$, assume the conditional distribution of $o_{t}$ is the normal distribution $N\left(w_{\gamma}^{T}\texttt{TF}_{\hat{\theta}}(H_{t}),\tilde{\sigma}^{2}\right)$, then we have
@@
    $\log\left(\frac{\mathbb{P}_{\gamma_{i}}(o_{t}|H_{t},\texttt{TF}_{\hat{\theta}})}{\mathbb{P}_{\gamma_{j}}\left(o_{t}|H_{t},\texttt{TF}_{\hat{\theta}}\right)}\right)$
    $=$ $\frac{\left(o_{t}-w_{\gamma_{j}}^{T}\texttt{TF}_{\hat{\theta}}(H_{t})\right)^{2}}{2\tilde{\sigma}^{2}}-\frac{\left(o_{t}-w_{\gamma_{i}}^{T}\texttt{TF}_{\hat{\theta}}(H_{t})\right)^{2}}{2\tilde{\sigma}^{2}}$
    $=$ $\frac{\left(w_{\gamma_{i}}^{T}\texttt{TF}_{\hat{\theta}}(H_{t})-w_{\gamma_{j}}^{T}\texttt{TF}_{\hat{\theta}}(H_{t})\right)o_{t}}{\tilde{\sigma}^{2}}+\frac{\left(w_{\gamma_{j}}^{T}\texttt{TF}_{\hat{\theta}}(H_{t})\right)^{2}-\left(w_{\gamma_{i}}^{T}\texttt{TF}_{\hat{\theta}}(H_{t})\right)^{2}}{2\tilde{\sigma}^{2}},$
    which is $D^{2}$-subgaussian with respect to $\gamma_{1}$, where $D$ is the bounded parameter of observation space and action space.
  • For dynamic pricing problem, given $\gamma,a_{1},o_{1},\ldots,a_{t}$, assume the conditional distribution of the demand $o_{t}$ is the normal distribution $\mathcal{N}\left(e_{\gamma}-f_{\gamma}a_{t},1\right)$, where $(e_{\gamma},f_{\gamma})$ is in the ball with center 0 and radius $\frac{D}{2}$, then we have
@@
    $\log\left(\frac{\mathbb{P}_{\gamma_{i}}(o_{t}|H_{t},\texttt{TF}_{\hat{\theta}})}{\mathbb{P}_{\gamma_{j}}\left(o_{t}|H_{t},\texttt{TF}_{\hat{\theta}}\right)}\right)$
    $=$ $\frac{\left(o_{t}-\left(e_{\gamma_{j}}-f_{\gamma_{j}}\texttt{TF}_{\hat{\theta}}(H_{t})\right)\right)^{2}}{2\tilde{\sigma}^{2}}-\frac{\left(o_{t}-\left(e_{\gamma_{i}}-f_{\gamma_{1}}\texttt{TF}_{\hat{\theta}}(H_{t})\right)\right)^{2}}{2\tilde{\sigma}^{2}}$
    $=$ $\frac{\left((f_{\gamma_{j}}-f_{\gamma_{i}})\texttt{TF}_{\hat{\theta}}(H_{t})+e_{\gamma_{i}}-e_{\gamma_{j}}\right)o_{\tau}}{\tilde{\sigma}^{2}}+\frac{\left(e_{\gamma_{j}}-f_{\gamma_{j}}\texttt{TF}_{\hat{\theta}}(H_{t})\right)^{2}-\left(e_{\gamma_{i}}-f_{\gamma_{i}}\texttt{TF}_{\hat{\theta}}(H_{t})\right)^{2}}{2\tilde{\sigma}^{2}},$
    which is $D^{2}+D$-subgaussian with respect to $\gamma_{1}$.
(2). To validate the condition that $\text{KL}\left(\mathcal{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\|\mathcal{P}_{\gamma_{j}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\right)\geq\Delta$, we illustrate with two examples: linear bandits and dynamic pricing. In these setting, we show that there exists a region in the neighborhood of $\texttt{Alg}^{*}(H_{t})$ where $\texttt{TF}_{\hat{\theta}}(H_{t})$ will satisfy this condition.
More specifically we show that this condition can be achieved when $\texttt{TF}_{\hat{\theta}}(H_{t})$ is defined as
$\texttt{TF}_{\hat{\theta}}(H_{t})=\left\{\begin{array}[]{lcl}\texttt{Alg}^{*}(H_{t}),&&\text{with probability $1-\epsilon$},\\ \bar{a},&&{\text{with probability $\epsilon$}},\\ \end{array}\right.$
for some $\bar{a}$.
  • For linear bandits problem,
@@
    $\text{KL}\left(\mathcal{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\|\mathcal{P}_{\gamma_{j}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\right)$
    $=$ $\epsilon\cdot\frac{\left(w_{\gamma_{i}}^{T}\bar{a}-w_{\gamma_{j}}^{T}\bar{a}\right)^{2}}{2\tilde{\sigma}^{2}}+(1-\epsilon)\cdot\frac{\left(w_{\gamma_{i}}^{T}\texttt{Alg}^{*}(H_{\tau})-w_{\gamma_{j}}^{T}\texttt{Alg}^{*}(H_{\tau})\right)^{2}}{2\tilde{\sigma}^{2}}.$
    Since the second term is always non-negative, one can choose some $\bar{a}$ which only depends on $w_{\gamma_{1}},\ldots,w_{\gamma_{k}}$ such that $\text{KL}\left(\mathcal{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\|\mathcal{P}_{\gamma_{j}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\right)\geq\epsilon\tilde{C}$ for some constant $\tilde{C}$, which also shows the trade-off between $\Delta$ and $\epsilon$ in Proposition 4.4.
  • For dynamic pricing problem,
@@
    $\text{KL}\left(\mathcal{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\|\mathcal{P}_{\gamma_{j}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\right)$
    $=$ $\epsilon\cdot\frac{\left((f_{\gamma_{j}}-f_{\gamma_{i}})\bar{a}+e_{\gamma_{i}}-e_{\gamma_{j}}\right)^{2}}{2\tilde{\sigma}^{2}}+(1-\epsilon)\cdot\frac{\left((f_{\gamma_{j}}-f_{\gamma_{i}})\texttt{Alg}^{*}(H_{\tau})+e_{\gamma_{i}}-e_{\gamma_{j}}\right)^{2}}{2\tilde{\sigma}^{2}}.$
    Then we can choose some $\bar{a}$ which only depends on $(e_{\gamma_{1}},f_{\gamma_{1}}),\ldots,(e_{\gamma_{k}},f_{\gamma_{k}})$ such that $\text{KL}\left(\mathcal{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\|\mathcal{P}_{\gamma_{j}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\right)\geq\epsilon\tilde{C}\triangleq\Delta$ for some constant $\tilde{C}$.

### v2

@@ Appendix D Proofs
For the dynamic pricing problem, we can construct a similar example. Suppose we have two environments without context $X_{t}$, denoted by $\gamma_{1},\gamma_{2}$. The demands $o_{t}$ of them are set to be $o_{t}=2-a_{t}+\epsilon_{t}$ and $o_{t}=\frac{4}{5}-\frac{1}{5}\cdot a_{t}+\epsilon_{t}$ respectively, where $\epsilon_{t}\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$ and $a_{t}$ is the price. Accordingly, the optimal actions are then $a^{(1)*}=1$ and $a^{(2)*}=2$.
Then if $\texttt{Alg}^{*}$ is the posterior averaging, we have
$a_{t}=\texttt{Alg}^{*}(H_{t})=\mathcal{P}_{\gamma}\left(\gamma_{1}|H_{t}\right)\cdot a^{(1)*}+\mathcal{P}_{\gamma}\left(\gamma_{2}|H_{t}\right)\cdot a^{(2)*},$
where the posterior distribution $\mathcal{P}_{\gamma}\left(\gamma_{i}|H_{t}\right)$ is given by
$\mathcal{P}_{\gamma}\left(\gamma_{i}|H_{t}\right)=\frac{\prod_{\tau=1}^{t-1}\mathbb{P}_{\gamma_{i}}(o_{\tau}|a_{\tau})}{\sum_{i^{\prime}=1}^{2}\prod_{\tau=1}^{t-1}\mathbb{P}_{\gamma_{i^{\prime}}}(o_{\tau}|a_{\tau})},\quad i=1,2.$
with $\mathbb{P}_{\gamma_{1}}(o_{\tau}|a_{\tau})$ as the normal distribution $\mathcal{N}\left(2-a_{t},1\right)$ and $\mathbb{P}_{\gamma_{2}}(o_{\tau}|a_{\tau})$ as $\mathcal{N}\left(\frac{4-a_{t}}{5},1\right)$.
Observe that $a_{1}=\frac{1}{2}a^{(1)*}+\frac{1}{2}a^{(2)*}=\frac{3}{2}$ satisfies $2-a_{1}=\frac{4-a_{1}}{5}$. Following a similar analysis as in the linear bandits example, we can conclude that $\mathcal{P}_{\gamma}\left(\gamma_{i}|H_{t}\right)=\frac{1}{2}$ for any $t$ and $i=1,2$, where $H_{t}$ can be generated by either $\gamma_{1}$ or $\gamma_{2}$. Therefore, $a_{t}=\frac{3}{2}$ for all $t\geq 1$ and the regret is
@@
$\text{Regret}(\texttt{Alg}^{*};\gamma_{i})=\mathbb{E}\left[\sum_{t=1}^{T}r(X_{t},a_{t}^{*})-r(X_{t},a_{t})\right]=\frac{1}{4}T\cdot\mathbb{I}\left\{\gamma_{i}=\gamma_{1}\right\}+\frac{1}{20}T\cdot\mathbb{I}\left\{\gamma_{i}=\gamma_{2}\right\},$
for $i=1,2$.
@@
We make the following additional assumptions:
  • There exists a constant $\bar{r}$ such that $\sup_{x\in\mathcal{X},a\in\mathcal{A},\gamma\in\Gamma}r(x,a;\gamma)\leq\bar{r}$, where we use $r(x,a;\gamma)$ to denote the reward function of environment $\gamma$.
  • There exists a constant $C_{r}$ such that $\texttt{Alg}^{*}$ satisfies
    $\mathbb{E}[r(X_{t},a^{*}_{t})-r(X_{t},\texttt{Alg}^{*}(H_{t}))|H_{t}]\leq C_{r}\sum_{\gamma_{i}\neq\gamma}\mathcal{P}(\gamma_{i}|H_{t}),$
    where the expectation on the left side is taken with respect to the possible randomness in $\texttt{Alg}^{*}$.
  • $\mathcal{P}_{\gamma}$ is a uniform distribution over $\Gamma$.
Proof.
We first compute a concentration inequality for $\sum_{\tau=1}^{t}\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|X_{\tau},a_{\tau})}{\mathbb{P}_{\gamma^{\prime}}(o_{\tau}|X_{\tau},a_{\tau})}\right)$. By the Bernstein-type concentration bound for a martingale difference sequence (Theorem 2.19 in [Wainwright, 2019]), under the given conditions, we have for any $\gamma^{\prime}\in\Gamma$, $t>0$ and $01>\delta>$ 0 with probability $1-\delta$,
@@
$\sum_{\tau=1}^{t}\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|X_{\tau},a_{\tau})}{\mathbb{P}_{\gamma^{\prime}}(o_{\tau}|X_{\tau},a_{\tau})}\right)-\mathbb{E}\left[\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|X_{\tau},a_{\tau})}{\mathbb{P}_{\gamma^{\prime}}(o_{\tau}|X_{\tau},a_{\tau})}\right)\right]\geq-\sqrt{2C_{\sigma^{2}}\Delta_{\text{Explore}}t\log(1/\delta)},$
where
@@
$\sum_{\tau=1}^{t}\mathbb{E}\left[\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|X_{\tau},a_{\tau})}{\mathbb{P}_{\gamma^{\prime}}(o_{\tau}|X_{\tau},a_{\tau})}\right)\right]=\sum_{\tau=1}^{t}\text{KL}\left(\mathbb{P}_{\gamma}(\cdot\big{|}X_{\tau},a_{\tau})\|\mathbb{P}_{\gamma^{\prime}}(\cdot\big{|}X_{\tau},a_{\tau})\right)\geq t\Delta_{\text{Explore}}.$
We think about two situations:
  • If $\sqrt{t\Delta_{\text{Explore}}}\geq 2\sqrt{2C_{\sigma^{2}}\log(1/\delta)}$, then
@@
    $\sum_{\tau=1}^{t}\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|X_{\tau},a_{\tau})}{\mathbb{P}_{\gamma^{\prime}}(o_{\tau}|X_{\tau},a_{\tau})}\right)$ $\geq t\Delta_{\text{Explore}}-\frac{1}{2}t\Delta_{\text{Explore}}$
    $=\frac{1}{2}t\Delta_{\text{Explore}}$
    $>\frac{1}{2}t\Delta_{\text{Explore}}-4C_{\sigma^{2}}\log(1/\delta).$
  • If $\sqrt{t\Delta_{\text{Explore}}}<2\sqrt{2C_{\sigma^{2}}\log(1/\delta)}$, then
    $\sum_{\tau=1}^{t}\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|X_{\tau},a_{\tau})}{\mathbb{P}_{\gamma^{\prime}}(o_{\tau}|X_{\tau},a_{\tau})}\right)$ $>t\Delta_{\text{Explore}}-4C_{\sigma^{2}}\log(1/\delta)$
    $\geq\frac{1}{2}t\Delta_{\text{Explore}}-4C_{\sigma^{2}}\log(1/\delta).$
Thus with probability $1-\delta$, we have
$\sum_{\tau=1}^{t}\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|X_{\tau},a_{\tau})}{\mathbb{P}_{\gamma^{\prime}}(o_{\tau}|X_{\tau},a_{\tau})}\right)>\frac{1}{2}t\Delta_{\text{Explore}}-4C_{\sigma^{2}}\log(1/\delta).$
Now we apply the union bound to $\gamma_{i}$ for $i=1,\ldots,k$ and for $t=1,\ldots,T$: with probability $1-1/T$, for all $i=1,\ldots,k$, and $t=1,\ldots,T$
@@
$\mathcal{P}(\gamma_{i}|H_{t})$ $\leq\mathcal{P}(\gamma|H_{t})\cdot\exp\left(-\frac{1}{2}t\Delta_{\text{Explore}}+4C_{\sigma^{2}}\log(1/\delta)\right).$
Now we are ready to prove the regret bound. We first decompose the regret:
$\text{Regret}(\texttt{TF}_{\hat{\theta}};\gamma)$ $=\sum_{t=1}^{T}\mathbb{E}[r(X_{t},\texttt{Alg}^{*}(H_{t}))-r(X_{t},\texttt{TF}_{\hat{\theta}}(H_{t}))]+\sum_{t=1}^{T}\mathbb{E}[r(X_{t},a^{*}_{t})-r(X_{t},\texttt{Alg}^{*}(H_{t}))]$
@@
$\leq\Delta_{\text{Exploit}}T+C_{r}\mathbb{E}\left[\sum_{t=1}^{T}\sum_{\gamma_{i}\neq\gamma}\mathcal{P}(\gamma_{i}|H_{t})\right].$
Since $\mathcal{P}(\gamma_{i}|H_{t})+\mathcal{P}(\gamma|H_{t})\leq 1$, we have with probability $1-1/T$, for all $i=1,\ldots,k$, and $t=1,\ldots,T$,
$\mathcal{P}(\gamma_{i}|H_{t})\leq\frac{1}{1+\exp\left(\frac{1}{2}t\Delta_{\text{Explore}}-4C_{\sigma^{2}}\log(kT^{2})\right)}.$
Let $t_{0}=\left\lceil{\frac{16C_{\sigma^{2}}\log(kT^{2})}{\Delta_{\text{Explore}}}}\right\rceil$, then we have with probability $1-1/T$, for all $i=1,\ldots,k$, and $t=1,\ldots,T$,
$\mathcal{P}(\gamma_{i}|H_{t})\leq\frac{1}{1+\exp\left(\frac{1}{2}t\Delta_{\text{Explore}}-4C_{\sigma^{2}}\log(kT^{2})\right)}<\exp\left(-\frac{1}{4}t\Delta_{\text{Explore}}\right).$
Thus, we can conclude that
$\mathbb{E}\left[\sum_{t=1}^{T}\sum_{\gamma_{i}\neq\gamma}\mathcal{P}(\gamma_{i}|H_{t})\right]$ $<1+\frac{16C_{\sigma^{2}}\bar{r}\log(kT^{2})}{\Delta_{\text{Explore}}}+k\mathbb{E}\left[\sum_{t=t_{0}}^{T}\exp\left(-\frac{1}{4}t\Delta_{\text{Explore}}\right)\right]$
$=O\left(\frac{\log k}{\Delta_{\text{Explore}}}\right)$
and finish the regret bound.
D.7 Example/Justification for the conditions in Proposition 4.4
We use a set of linear bandits problems as a an example to illustrate Proposition 4.4. Consider a set of linear bandits problems with $o_{t}=R_{t}=w^{\top}a_{t}+\epsilon_{t}$, where $\epsilon_{t}$ is i.i.d. from a standard Normal distribution. Recall the environment parameter $\gamma=w$ here. We assume it has dimension $d$ and both $\mathcal{A}$ and $\Gamma$ are bounded with respect to the $L2$ norm. Further, we assume $\texttt{Alg}^{*}(H_{t})$ is posterior averaging and for each time $t$,
$\texttt{TF}_{\hat{\theta}}(H_{t})=\texttt{Alg}^{*}(H_{t})+\Delta_{t},$
where $\Delta_{t}$ follows a uniform distribution in $[-1,1]^{d}$, i.i.d. across time and independent of $H_{t}$, $a_{t}$, and $\gamma$. This follows the numerical observations in Figure 3.
In this setting, we verify the conditions in Proposition 4.4. To simplify notations, we denote $\tilde{a}^{*}_{t}=\texttt{Alg}^{*}(H_{t})$.
First condition. Since
$\mathbb{E}[r(X_{t},\texttt{Alg}^{*}(H_{t}))-r(X_{t},\texttt{TF}_{\hat{\theta}}(H_{t}))|H_{t}]=\mathbb{E}[\gamma^{\top}\tilde{a}^{*}_{t}-\gamma^{\top}\tilde{a}^{*}_{t}-\gamma^{\top}\Delta_{t}|H_{t}]=0,$
we have $\Delta_{\text{Exploit}}=0.$
Second condition. Since $\epsilon_{t}$ is i.i.d. from a standard Normal distribution, we have for $o_{t}=\gamma^{\top}a_{t}+\epsilon_{t}$ and $a_{t}=\texttt{TF}_{\hat{\theta}}(H_{t})$,
$\log\left(\frac{\mathbb{P}_{\gamma}(o_{t}|X_{t},a_{t})}{\mathbb{P}_{\gamma^{\prime}}(o_{t}|X_{t},a_{t})}\right)$ $=\frac{(o_{t}-\gamma^{{}^{\prime}\top}a_{t})^{2}-(o_{t}-\gamma^{\top}a_{t})^{2}}{2}$
$=\frac{((\gamma-\gamma^{\prime})^{\top}a_{t})^{2}}{2}-(\gamma-\gamma^{\prime})^{\top}a_{t}\epsilon_{t}.$
Thus, the KL divergence, which is the expectation of the above term with respect to $\epsilon_{t}$ and $\Delta_{t}$ (which are independent of $H_{t}$), is
$\mathbb{E}\left[\frac{((\gamma-\gamma^{\prime})^{\top}(\tilde{a}^{*}_{t}+\Delta_{t}))^{2}}{2}\bigg{|}H_{t}\right]\geq\frac{1}{6}\|\gamma-\gamma^{\prime}\|_{2}^{2}.$
And further, for the “noise” term $(\gamma-\gamma^{\prime})^{\top}a_{t}\epsilon_{t}$, since its variance is bounded by $O(\|\gamma-\gamma^{\prime}\|_{2}^{2})$ (due to the assumptions on the boundedness of $a_{t}$ and that $\epsilon_{t}$ follows a standard Normal distribution), we can set $\Delta_{\text{Explore}}=\min_{\gamma^{\prime}\in\Gamma}\|\gamma-\gamma^{\prime}\|_{2}^{2}.$
Further, since
$\mathbb{E}[r(X_{t},a^{*}_{t})-r(X_{t},\texttt{Alg}^{*}(H_{t}))|H_{t}]\leq\bar{r}\sum_{\gamma_{i}\neq\gamma}\mathcal{P}(\gamma_{i}|H_{t}),$
we can set $C_{r}=\bar{r}$.
And thus, we can conclude the regret bound is $\text{Regret}(\texttt{TF}_{\hat{\theta}},\gamma)=O\left(\frac{\log k}{\min_{\gamma^{\prime}\in\Gamma}\|\gamma-\gamma^{\prime}\|_{2}^{2}}\right)$.